The balanced combustion reaction for C6H6 is
2C6H6(l)+15O2(g)⟶12CO2(g)+6H2O(l)+6542 kJ

If 5.500 g C6H6 is burned and the heat produced from the burning is added to 5691 g of water at 21 ∘ C, what is the final temperature of the water?
1
Expert's answer
2019-10-21T07:16:03-0400

n(C6H6) = 5.500/78 = 0.0705 mol

E = (6542*0.0705)/2 = 230.60 kJ

C(H2O) = 4.183 kJ/(kg*K)

T = 21+230.60/(4.183*5.691) =30.69 C